# Dimensionality increase for error correction in the interaction between information space and the physical world

### v1

@@
Abstract. The evolution of human intelligence led to the huge amount of data in the information space. Accessing and processing this data helps in finding solutions to applied problems based on finite-dimensional models. We argue, that formally, such a mathematical model can be embedded into a higher-dimensional model inside of which a desired solution will exist.
@@ 1. Introduction
There is a vast amount of literature on the computational and mathematical methods of learning (e.g. the monograph [11]), the interaction between a computer and a human, and theoretical aspects of information engineering related to signal processing or control theory. Geometric methods, especially those based on Riemannian geometry, have been extensively used in the study of networks, information transmission, and data analysis: see e.g. discussion and references in [2], [3].
@@
Predicting, analyzing and correcting errors (or malfunctions) is an essential part of data analysis and engineering. To give some specific examples, we mention the paper [9] that discusses the contributing factors to errors in machining and presents theoretical and quantitative analysis of the model that leads to improved accuracy. The word ”error”, in the most general sense, can be understood as an issue to be fixed or a problem to be addressed. In the abstract sense, it can be parameters or data that need to be modified. The paper [10] reviews the issues that occur in large networks, such as congestion, delays or packet loss, and possible approaches to gaining control over these problems. The paper [4] addresses the problem of low power in radio frequency energy harvesting systems. Mathematical representations or models of real life systems often serve as useful approximations. For instance, in [12], the model for an electrical vehicle system allows to compare and evaluate the system responses.
@@ 2. Main theorem and proof
where $U$, $V$ are (any) submanifolds of W, and $C$, $C_{1}$,…, $C_{k}$, for a positive integer k, are (arbitrary) submanifolds of E. These maps include, in particular, the maps that represent direct human-to-human interactions, or the human-to-human encoded communication of any form, or a person putting data into E or retrieving information from E.
@@
Theorem 1. Let $W$, $E$ be as defined above. Let $a<b$ be real numbers and let $I=[a,b]$. Let $n>1$ be the dimension of $S$, the submanifold of $W$ that models the physical world. Let $A$ be a submanifold of $S$. Assume the manifolds $S$, $A$ are compact, with or without boundary. Let $B$ be a finite-dimensional submanifold of $W$, diffeomorphic to $A$, such that $A\cap B=\emptyset$. Then there exist a submanifold $X$ of $W$ of dimension $n\leq\dim X\leq 2n+1$ and a smooth family of submanifolds $Y(t)$, $t\in I$, of $X$ such that $Y(a)=A$ and $Y(b)=B$.

### v2

@@
Abstract.
The evolution of human intelligence led to the huge amount of data in the information space. Accessing and processing this data helps in finding solutions to applied problems based on finite-dimensional models. We argue, that formally, such a mathematical model can be embedded into a higher-dimensional model inside of which a desired solution will exist. In our model, the physical world and the information space are submanifolds of infinite-dimensional Hilbert spaces, and the processes, including information transmission, are maps between the submanifolds of the physical world or of the information space. We discuss how our perspective fits in the context of existing literature. Our theorem states that a submanifold in the parameter space of the physical world can be deformed to a target submanifold outside that space, with an appropriate count of the deformation parameters. We interpret this assertion as an existence result for a class of problems and we discuss further steps.
@@ 1. Introduction
There is a vast amount of literature on the computational and mathematical methods of learning (e.g. the monograph [25]), the interaction between a computer and a human, and theoretical aspects of information engineering related to signal processing or control theory. Geometric methods, especially those based on Riemannian geometry, have been extensively used in the study of networks, information transmission, and data analysis. We note in particular the fundamental works of Amari, Belkin and Niyogi, and a recent series of papers by Fefferman and collaborators including [9]. More discussion and references can be found in [3], [4].
@@
In this paper, we state that the physical world is a submanifold of a manifold $W$, and the information space is a submanifold of a manifold $E$. We discuss the interactions (mappings) between the two. This approach resonates with a number of other perspectives such as
  • Gärdenfors’ conceptual spaces, with his emphasis on geometric structures [11], see also a more practical development in this general direction e.g. in [26],
  • digital twins (a general perspective in [22], and an example of application of digital twins to a practical problem in medicine in [17]),
  • world/foundation models [6], [15],
  • generally, mathematical modeling (see [14] for a discussion of how mathematical modeling differs from digital twins).
Predicting, analyzing and correcting errors (or malfunctions) is an essential part of data analysis and engineering. There is an extensive theory or error correction and fault tolerant computation. This massive variety of methods originated from the necessity to deal with the noise and the accumulation of errors that occur as the data packets are transmitted through various types of channels. The word ”error”, in the most general sense, can be understood as an issue to be fixed or a problem to be addressed. In the abstract sense, it can be parameters or data that need to be modified. Broadly, one can address the question of faulty signal propagation in a network, regardless of the type of the network. It could be a mathematical model for a network of neurons or nerves in a living organism, or for a societal network. In the daily functioning of human society, there are quantifiable errors in the information flow that may cause operational issues. For example, an office assistant may accidentally communicate a wrong meeting room number (a simple email typo), or a post office worker may place an envelope in a wrong mail box (as a result of being momentarily distracted).
There is extensive literature on error correction, including the foundational Shannon’s paper [23] and von Neumann’s paper [27]. While it is a difficult task to give a proper literature overview, we will mention a few other works. There has been a lof of effort to construct error correcting codes [10], [18]. Quantum error correction, with references, is discussed in [12] and Chapter 10 of [20].
We also include a few examples of works that address specific real life systems. The paper [21] discusses the contributing factors to errors in machining and presents theoretical and quantitative analysis of the model that leads to improved accuracy. The paper [24] reviews the issues that occur in large networks, such as congestion, delays or packet loss, and possible approaches to gaining control over these problems. The paper [7] addresses the problem of low power in radio frequency energy harvesting systems. Mathematical representations or models of real life systems often serve as useful approximations. For instance, in [29], the model for an electrical vehicle system allows to compare and evaluate the system responses.
@@ 2. Main theorem and proof
where $U$, $V$ are (any) submanifolds of W, and $C$, $C_{1}$,…, $C_{k}$, for a positive integer k, are (arbitrary) submanifolds of E. These maps include, in particular, the maps that represent direct human-to-human interactions, or the human-to-human encoded communication of any form, or a person putting data into E or retrieving information from E. This is consistent with regarding an information transfer as an embedding of manifolds (see e.g. the discussion in [3]).
@@
Example. In this example, we will provide a concrete example of an element of G and how it contributes to error correction. More examples and discussion are in Section 3 below.
Suppose Alice has to upload her Math 101 homework to the course site. She noted that assignment 1 is due October 1 (let’s say, this date is $d$, on the $d$-axis in a copy of $\mathbb{R}$ embedded in the information space), Alice retrieved this information from $E$ (by checking the homework sheet) and stored the value of $d$ (in her memory cells, or in her calendar). However, being in a rush, Alice missed that the deadline is 3:00 pm on October 1. On October 1, Alice attempted to upload her file homework1.pdf to the course site. This is a map $f_{1}:U\to V$ of submanifolds of $W$, that induces a map $f_{2}:V\to C_{1}\subset E$, the deadline check map $f_{3}:C_{1}\to C_{2}$ in $E$ that indicates that the submission arrived after the 3:00 pm deadline, and the resulting map $f_{4}:C_{2}\to W$ that produces a message to Alice in $W$ that the submission is closed and her paper can not be uploaded. The element of $G$ is $f:U\to W$ defined by the composition $f=f_{4}\circ f_{3}\circ f_{2}\circ f_{1}$. A way to fix the issue, for the next assignment submission, is to increase the number of parameters, and for Alice to retrieve and store two parameters instead of one: $d$, the date, and $t$, the time (in this case, 15:00 hours).
@@
Theorem 1. Let $W$, $E$ be as defined above. Let $a<b$ be real numbers and let $I=[a,b]$. Let $n>1$ be the dimension of $S$, the submanifold of $W$ that models the physical world. Let $A$ be a submanifold of $S$. Assume the manifolds $S$, AA are compact, with or without boundary. Let $B$ be a finite-dimensional submanifold of $W$, diffeomorphic to $A$, such that $A\cap B=\emptyset$. Then there exist a submanifold $X$ of $W$ of dimension $n\leq\dim X\leq 2n+1$ and a smooth family of submanifolds $Y(t)$, $t\in I$, of $X$ such that $Y(a)=A$ and $Y(b)=B$.
@@ 4. Conclusions
4.2. Future work and limitations.
This paper is an observation that ramified from [2], [3], [4]. The general line of thought (that in some sense has roots in [27], [28]) is to develop the mathematical methods allowing to deal with the issues that occur with electrical signal transmission in human body. The question addressed in this note is about solvability of the problem, and the answer is an existence statement for a solution (Theorem 1). This answer is not constructive, as it is only an assertion that a solution exists, and this can be regarded as a limitation. Another limitation is the setup of the mathematical model (in particular, the requirement that $B$ is diffeomorphic to $A$, even though we do not require that $B$ is a subset of $S$, which allows for some freedom, as it means that $B$ describes a configuration that is currently not realized in the physical world).
It is known that a number of issues with the human body (for example, certain instances of chronic pain) result from a malfunction in neurons, although there is no complete knowledge of these mechanisms or of the ways to repair it. A suitable direction for future work is towards a mathematical model that addresses the geometry of this situation and the appropriate choice of the parameter space. To handle the uncertainty, one possible approach is to use fuzzy manifolds, or a suitable generalization of this concept. This should work well with hidden Markov processes or quantum hidden Markov processes. Quantumness is a separate issue that needs to be addressed, as the signal propagation in neural networks in human body seems to be a ”partially classical and partially quantum” process. Replacing a probability measure $\rho$ on a manifold $M$ with a smooth membership function $\mu:M\to[0,1]$ would change our interpretation of $M$. In a simple example when $M$ is the open unit disk in the $xy$-plane and $\rho=\frac{1}{\pi}dA$, where $dA$ is the Lebesgue measure, there are many constant membership functions: $\mu_{c}:M\to[0,1]$, $\mu_{c}(x,y)\equiv c$, where $c$ is a constant between 0 and $1$. The function $\mu_{1}$ corresponds to the case when for every point $(x,y)$ such that $x^{2}+y^{2}<1$ we have $(x,y)\in M$ in the fuzzy set $(M,\mu_{1})$. The function $\mu_{0}$ corresponds to the case when for every point $(x,y)$ such that $x^{2}+y^{2}<1$, $(x,y)\notin M$ in the fuzzy set $(M,\mu_{0})$. Working with fuzzy manifolds is difficult, as it very quickly becomes clear already at the level of fuzzy topological spaces [5]. But, intuitively, this can be a right mathematical concept to support the study of quantitative processes that happen for an unknown reason (although the reason can possibly be identified in the correctly chosen parameter space).
Acknowledgment
T.B. thanks the anonymous reviewers for their valuable comments and suggestions.